# Probing the Origins I: Generalised Additive Model inference of birth radii for Milky Way stars in the solar vicinitythanks: Catalogue reference

### v1

@@ Abstract
Context. As stars traverse the Galaxy, interactions with structures like the bar and spiral arms can alter their orbits, leading to either “churning”, where changes in angular momentum shift their guiding radii, or “blurring”, where angular momentum is preserved. Churning is what is commonly known as radial migration.
Aims. Here, we probe the orbital characteristics of a diverse set of stars in the thin disc observed by the Gaia-ESO survey. We aim to discern whether their orbits are predominantly influenced by churning or if they kept their orbital birth radii (i.e. were blurred or remained undisturbed).
Methods. We employ a Generalised Additive Model (GAM) to address the limitations inherent in radial metallicity gradients predicted by chemical evolution models, thereby facilitating the estimation of birth radii for the thin disc stars in our sample based on their ages and chemical composition. We then juxtapose the birth radius predictions derived from the GAM with the calculated guiding radii, among other dynamic parameters. This comparison is performed within distinct groups of our dataset, categorised through Hierarchical Clustering (HC) based on 21 chemical abundances spanning 18 species.
@@ 1 Introduction
The Milky Way (MW) is a complex combination of stars, gas, dust, and dark matter (e.g. Binney & Vasiliev, 2024, and references therein). Detailed studies of the MW are important to contextualise its role as the cradle of our Solar System (Gonzalez et al., 2001; Stojković et al., 2019) and as a fundamental archetype for unravelling galactic structures and evolutionary processes, which makes the MW a reference for studying other galaxies (e.g. Kobayashi & Taylor, 2023, and references therein).
Our Galaxy is a barred spiral with several substructures: the thin and thick discs, the halo, the (box/peanut-shaped) bulge, the spiral arms, and the bar, all interacting intricately (Bland-Hawthorn & Gerhard, 2016). As new stars form and evolve, they are influenced by these structures and their gravitational tugs, which can perturb their orbits (e.g. Sellwood & Binney, 2002; Lépine et al., 2003; Roškar et al., 2008; Brunetti et al., 2011; Martínez-Bautista et al., 2021; Carr et al., 2022; Lu et al., 2022; Fujimoto et al., 2023; Iles et al., 2024; Nepal et al., 2024). Some stars may have been formed approximately in their current Galactocentric orbital radii, while others may have migrated, disturbed by the complex interplay of forces within the MW. Additionally, there have been evidence that satellite galaxies could also induce radial mixing in the outer disc of MW-like galaxies (e.g. Quillen et al., 2009). By studying the stars in the solar vicinity, we can learn about the various Galactic stellar populations and gain insight into their histories and origins.
In this context, the dynamic process most commonly referred to as radial migration is the one known as churning, which involves changes in a star’s angular momentum, causing a shift in its guiding radius without necessarily affecting its orbital eccentricity. A related process, often discussed alongside churning, is blurring, where stars undergo epicyclic motion around their guiding radii, maintaining their angular momentum, and typically orbiting at high eccentricities (see, for instance, Sellwood, 2014; Halle et al., 2015; Frankel et al., 2020; Wozniak, 2020, and references therein).
@@
Indeed, in a previous work (Dantas et al., 2023), we explored a set of super-metal-rich stars currently inhabiting the solar vicinity that were observed by the Gaia-ESO public spectroscopic survey (Gilmore et al., 2012, 2022; Randich et al., 2013, 2022). These stars are old (typical median ages of $\sim$ 8 Gyr), have near-circular orbits (median eccentricities around 0.2), and can reach maximum Galactic heights between 0.5-1.5 kpc, all characteristics of stars with perturbed orbits due to the interactions with major Galactic structures, meaning the bar and/or spiral arms. In addition to migrating metal-rich stars, there is probably a fraction of metal-poor thin-disc stars that are also migrators but originate from the outer regions of the disc (Haywood, 2008).
@@
To try to quantify how much a star has been dislocated from its original birth radius, it is usually necessary to make assumptions about how the radial metallicity gradient of the interstellar medium changed with time (see, for instance, Frankel et al., 2018; Minchev et al., 2018; Feltzing et al., 2020; Ratcliffe et al., 2023). However, how the metallicity gradient changes with time is itself affected by the radial motions of gas and stars (see, for instance, Kubryk et al., 2015a, b). Observational constraints on the temporal variations of the metallicity gradient are important ingredients for models of Galactic chemical evolution. Studies in that direction have been done with samples of stars (e.g. Xiang et al., 2015; Anders et al., 2017; Willett et al., 2023) and open clusters (Chen & Zhao, 2020; Spina et al., 2021; Magrini et al., 2023) with precise ages. Young components of the Galaxy, such as Cepheids and some open clusters, are very useful, as they hardly have had enough time to interact with the Galaxy’s structures and, therefore, have probably not (yet) been subject to relocation (e.g. da Silva et al., 2023; Viscasillas Vázquez et al., 2023).
@@
Such chemical evolution models play a crucial role by mapping the spatial and temporal evolution of the MW. They are intrinsically connected to the idea of the “inside-out” formation and growth of the Galaxy. As per this conceptual framework, the initial phase of Galactic evolution entails the progressive consolidation of the central region, where star formation is intense and fast, quickly resulting in material of high metallicity. The outer regions of the disc are progressively consolidated, but with an outward gradient in star formation intensity, resulting in temporal differences in their chemical enrichment (e.g. Chiappini et al., 2001; Kobayashi et al., 2006; Magrini et al., 2009; Bergemann et al., 2014; Andrews et al., 2017; Schönrich & McMillan, 2017; Hu et al., 2023; Magrini et al., 2023). Additionally, there is a myriad of evidence suggesting that the inside-out formation scenario also happens in other galaxies, not necessarily only in spiral ones (e.g. Bezanson et al., 2009; Avila-Reese et al., 2023).
@@
Previous studies have estimated stellar birth radii ($R_{\rm b}$) using various datasets and methodologies. For instance, Feltzing et al. (2020) and Ratcliffe et al. (2023) used data from the Apache Point Observatory Galactic Evolution Experiment (APOGEE; Majewski et al., 2017). While Feltzing et al. (2020) relied on various distinct chemical enrichment models to derive $R_{\rm b}$, aiming to constrain and broaden the range of plausible $R_{\rm b}$, Ratcliffe et al. (2023) employed the empirical approach developed in Lu et al. (2024). Similarly, Minchev et al. (2018) adopted a semi-empirical method, applying it to a sample of stars observed with the High Accuracy Radial velocity Planet Searcher instrument (HARPS; Mayor et al., 2003). These diverse approaches are valuable, as they provide important benchmarks for assessing and comparing new techniques and methodologies.
@@
In this paper, we propose an alternative approach for deriving $R_{\rm b}$, that serves as an additional framework to complement and compare with previous studies. Specifically, we employ Generalised Additive Models (GAMs; introduced by Hastie & Tibshirani, 1990) to expand theoretical chemical evolution models for the thin disc (in this case, those of Magrini et al. 2009) of the MW (which are characterised by widely spaced bins in age and radius) to investigate and differentiate the motion of thin disc stars (i.e. churn and blur/lack of interaction). Our GAM model incorporates two independent variables consisting of chemical abundances ([Fe/H] or [Mg/H]; the analysis with [Mg/H] is kept in the Appendix) and $t$, the age of the Universe at the time the star was formed (roughly 13.8 Gyr [^1] We adopt the estimated age of the Universe as 13.8 Gyr (Planck Collaboration et al., 2020). minus the current estimated age of a star); while the Galactocentric distance ($R$) serves as the response variable (i.e. the output of our regression model). Using GAM, we can effectively model the complex relationship between the chemical abundances and $t$ (independent variables), and $R$ without imposing rigid assumptions regarding the functional forms of these relationships.
@@
In contrast, the thick disc – despite its distinct [$\alpha$/Fe] enhancements, older stellar populations, and kinematically hot nature (see e.g. Soubiran et al., 2003; Bensby et al., 2011; Recio-Blanco et al., 2014; Bland-Hawthorn & Gerhard, 2016) – presents greater complexities that remain less well understood. Structures such as the halo and bulge pose even greater challenges for chemical evolution models, as their diverse formation mechanisms result in extreme variations in metallicity and dynamics. For instance, the halo reflects the earliest stages of Galactic chemical evolution and accretion events, being predominantly metal-poor (see e.g. Gonzalez-Jara et al., 2024). In contrast, the bulge, characterised by its old stellar populations and strikingly wide metallicity variations ($-1.5\lesssim\rm{[Fe/H]}\lesssim+0.5$; see Barbuy et al., 2018, for a comprehensive review), encapsulates the star formation history of the Galactic centre. Despite its importance, providing a self-consistent empirical chemo-dynamical model for the Galactic bulge remains a significant challenge due to its complexity (e.g. Grieco et al., 2012; Barbuy et al., 2018).
Insights from numerical simulations, particularly regarding metallicity gradients and dynamical evolution across these Galactic components, provide valuable context for understanding their formation and evolution (e.g. Grand et al., 2017; Tissera et al., 2022; Jara-Ferreira et al., 2024). However, these theoretical predictions must be confronted with observations to ensure consistency with the MW’s history and evolution. Consequently, while the thin disc offers a tractable and data-rich framework for advancing our understanding of chemical evolution, expanding these models to incorporate other Galactic components – guided by both simulations and observational constraints – remains an essential and ongoing objective (see e.g. Minchev et al., 2013, for a study combining chemical enrichment models of the MW’s disc with numerical simulations of galactic discs).
This work forms the foundation of a broader investigation. Here, we focus on analysing results derived from the GAM, providing an in-depth exploration of our observed Gaia-ESO star sample. Our goal is to examine other aspects of these findings in future studies to better understand the causes and implications of the stellar motion trends revealed here.
This paper is organised as follows: Section 2 details the dataset and methodologies used in this study, providing the foundation for the subsequent analysis. The heart of the paper lies in Section 3, where we delve into the analysis and discuss the implications of our findings. Concluding the paper, Section 4 synthesises these insights, summarising the key takeaways and the conclusions drawn from this phase of our investigation.
@@ 2.1 Data description
Figure 2: Scatter plot displaying the relationship between metallicity, [Fe/H], and stellar ages, $t_{\star}$, using unidam estimates for the entire sample. Data points are categorised by Hierarchical Clustering (HC) groups. On top, 1D-Gaussian kernel densities show age distribution; on the right, a similar view presents [Fe/H] distribution for each HC group. The window width for the kernel density estimation is computed using Scott’s method, with a default bandwidth adjustment of 1, which is the standard configuration of the seaborn package in python (Waskom, 2021); this configuration is consistently applied to all figures with Gaussian kernel densities throughout this paper. The data points and densities are colour-coded and stratified according to the different HC groups, similarly to Fig. 1.
@@
2.2 The Magrini et al.’s Galactic chemical evolution models
@@
In Fig. 3 the radial profiles of the Magrini et al. (2009) models are shown. It is noticeable that the bottom panel, depicting [Fe/H], shows the smoothest curves (potentially the best estimated of the profiles), whereas all the other ones have constant or near-constant abundances after a certain $R$. This is an important feature to have in mind since this lack of smoothness can be detrimental to the implementation of the GAM. This will be further discussed in Sect. 2.3. Moreover, there is a clear pattern of strong positive or negative correlations between the abundances in the models, as evidenced by the heatmap in Fig. 4.
@@
Figure 3: We present the MW’s chemical enrichment models as described by Magrini et al. (2009), illustrating the relative abundances of key elements in comparison to iron (Fe) across various Galactocentric radii ($R$). The elements, arranged sequentially by atomic number relative to Fe, include [O/Fe] at the top panel, followed by [Mg/Fe], [Si/Fe], [Ca/Fe], and culminating with [Fe/H] on the bottom panel. The colour gradation, ranging from a lively orange to black, represents the ages of the Universe at the time of star formation within the MW, starting at $t$=2.1 Gyr and extending to $t$=13.7 Gyr. This visual spectrum effectively portrays the chronological progression of the Galaxy’s chemical composition. The legend, for clearer visualisation, is incorporated within the [Ca/Fe] subplot.
Figure 4: Heatmap visualisation of the correlation matrix showing the pairwise correlation coefficients between elemental abundances relative to iron ([O/Fe], [Mg/Fe], [Si/Fe], [Ca/Fe]) and iron abundance relative to hydrogen ([Fe/H]). Orange indicates a strong positive correlation, while brown signifies a strong negative correlation, demonstrating the tight coupling between these elemental abundances in the context of Magrini et al.’s model. The correlation matrix was estimated via Spearman’s rank correlation coefficient (Spearman, 1904).
@@ 2.3 Generalised Additive Models
2.3.1 Why Generalised Additive Models?
@@
Generalised Additive Models (GAMs; introduced by Hastie & Tibshirani, 1990), an extension of GLMs, offer even greater flexibility by allowing for non-linear relationships between predictors and the response variable, using smooth functions. Despite their potential, to our knowledge, GAMs have been previously used fewer times in astrophysics, such as in photometric redshift estimation (Beck et al., 2017), in the analysis of ionising photon escape fractions from dark haloes (Hattab et al., 2019), and more recently in the examination of the relationship between galaxy structure and star formation rate (Stephenson et al., 2024).
In recent years, astrophysics seems to have skipped over these “intermediate” techniques, with many researchers jumping directly from classical methods to highly complex approaches like Gaussian Processes and Deep Learning. While these advanced techniques are undoubtedly valuable, methods such as GLMs and GAMs offer elegant and interpretable solutions that are more straightforward to implement and understand. This accessibility makes them highly valuable in a range of disciplines, shown by its use in ecology (e.g. Simpson, 2018; Pedersen et al., 2019; Kosicki, 2020), linguistics (e.g. Wieling et al., 2016; Tomaschek et al., 2018), medicine (e.g. de Souza & Berger, 2021), to mention a few.
2.3.2 The general form of a GAM

### v2

@@ Abstract
Context. As stars traverse the Galaxy, interactions with structures such as the bar and spiral arms can alter their orbits, leading either to ‘churning’, where changes in angular momentum shift their guiding radii, or ‘blurring’, where angular momentum is preserved. Churning is what is commonly known as radial migration.
Aims. Here, we probe the orbital characteristics of a diverse set of stars in the thin disc observed by the Gaia-ESO survey. We aim to discern whether their orbits are predominantly influenced by churning or if they keep their orbital birth radii (i.e. were blurred or remained undisturbed).
Methods. We employed a generalised additive model (GAM) to address the limitations inherent in radial metallicity gradients predicted by chemical evolution models, thereby facilitating estimation of the birth radii for the thin disc stars in our sample based on their age and chemical composition. We then juxtaposed the birth radius predictions derived from the GAM with the calculated guiding radii, among other dynamic parameters. This comparison was performed within distinct groups of our dataset, categorised through hierarchical clustering (HC) based on 21 chemical abundances spanning 18 species.
@@ 1 Introduction
The Milky Way (MW) is a complex combination of stars, gas, dust, and dark matter (e.g. Binney & Vasiliev, 2024, and references therein). Detailed studies of the MW are important for contextualising its role as the cradle of our Solar System (Gonzalez et al., 2001; Stojković et al., 2019) and as a fundamental archetype for unravelling galactic structures and evolutionary processes, which makes the MW a reference for studying other galaxies (e.g. Kobayashi & Taylor, 2023, and references therein).
Our Galaxy is a barred spiral with several substructures – thin and thick discs, the halo, the (box/peanut-shaped) bulge, spiral arms, and the bar – all interacting intricately (Bland-Hawthorn & Gerhard, 2016). As new stars form and evolve, they are influenced by these structures and their gravitational tugs, which can perturb their orbits (e.g. Sellwood & Binney, 2002; Lépine et al., 2003; Roškar et al., 2008; Brunetti et al., 2011; Martínez-Bautista et al., 2021; Carr et al., 2022; Lu et al., 2022; Fujimoto et al., 2023; Iles et al., 2024; Nepal et al., 2024). Some stars may have been formed near their current Galactocentric orbital radii, while others may have migrated, a result of being disturbed by the complex interplay of forces within the MW. Additionally, there is evidence that satellite galaxies could also induce radial mixing in the outer disc of MW-like galaxies (e.g. Quillen et al., 2009). By studying the stars in the solar vicinity, one can learn about the various Galactic stellar populations and gain insight into their histories and origins.
In this context, the dynamic process most commonly referred to as radial migration is known as churning. It involves changes in a star’s angular momentum that cause a shift in its guiding radius without necessarily affecting its orbital eccentricity. A related process often discussed alongside churning is blurring, which is a process where stars undergo epicyclic motion around their guiding radii, maintaining their angular momentum, and typically orbiting at high eccentricities (see, for instance, Sellwood, 2014; Halle et al., 2015; Frankel et al., 2020; Wozniak, 2020, and references therein).
@@
Indeed, in a previous work (Dantas et al., 2023), we explored a set of super-metal-rich stars currently inhabiting the solar vicinity that were observed by the Gaia-ESO public spectroscopic survey (Gilmore et al., 2012, 2022; Randich et al., 2013, 2022). These stars are old (typical median ages of $\sim$ 8 Gyr), have near-circular orbits (median eccentricities around 0.2), and can reach maximum Galactic heights between 0.5-1.5 kpc. These are all characteristics of stars with perturbed orbits due to the interactions with major Galactic structures, that is, the bar and/or spiral arms. In addition to migrating metal-rich stars, there is probably a fraction of metal-poor thin-disc stars that are also migrators but originate from the outer regions of the disc (Haywood, 2008).
@@
To try to quantify how much a star has been dislocated from its original birth radius, it is usually necessary to make assumptions about how the radial metallicity gradient of the interstellar medium changed with time (see, for instance, Frankel et al., 2018; Minchev et al., 2018; Feltzing et al., 2020; Ratcliffe et al., 2023). However, how the metallicity gradient changes with time is itself affected by the radial motions of gas and stars (see, for instance, Kubryk et al., 2015a, b). Observational constraints on the temporal variations of the metallicity gradient are important ingredients for models of Galactic chemical evolution. Studies in that direction have been done with samples of stars (e.g. Xiang et al., 2015; Anders et al., 2017; Willett et al., 2023) and open clusters (Chen & Zhao, 2020; Spina et al., 2021; Magrini et al., 2023) with precise ages. Young components of the Galaxy, such as Cepheids and some open clusters, are very useful, as they have hardly had enough time to interact with the Galaxy’s structures and therefore have probably not (yet) been subject to relocation (e.g. da Silva et al., 2023; Viscasillas Vázquez et al., 2023).
@@
Such chemical evolution models play a crucial role by mapping the spatial and temporal evolution of the MW. They are intrinsically connected to the idea of the ‘inside-out’ formation and growth of the Galaxy. As per this conceptual framework, the initial phase of Galactic evolution entails the progressive consolidation of the central region, where star formation is intense and fast and quickly results in material of high metallicity. The outer regions of the disc are progressively consolidated but with an outward gradient in star formation intensity, resulting in temporal differences in their chemical enrichment (e.g. Chiappini et al., 2001; Kobayashi et al., 2006; Magrini et al., 2009; Bergemann et al., 2014; Andrews et al., 2017; Schönrich & McMillan, 2017; Hu et al., 2023; Magrini et al., 2023). Additionally, there is myriad evidence suggesting that the inside-out formation scenario also happens in other galaxies, not necessarily only in spiral ones (e.g. Bezanson et al., 2009; Avila-Reese et al., 2023).
@@
Previous studies have estimated stellar birth radii ($R_{\rm b}$) using various datasets and methodologies. For instance, Feltzing et al. (2020) and Ratcliffe et al. (2023) used data from the Apache Point Observatory Galactic Evolution Experiment (APOGEE; Majewski et al., 2017), while Feltzing et al. (2020) relied on various distinct chemical enrichment models to derive $R_{\rm b}$, aiming to constrain and broaden the range of plausible $R_{\rm b}$. Ratcliffe et al. (2023) employed the empirical approach developed in Lu et al. (2024). Similarly, Minchev et al. (2018) adopted a semi-empirical method, applying it to a sample of stars observed with the High Accuracy Radial velocity Planet Searcher instrument (HARPS; Mayor et al., 2003). These diverse approaches are valuable, as they provide important benchmarks for assessing and comparing new techniques and methodologies.
@@
In this paper, we propose an alternative approach for deriving $R_{\rm b}$ that serves as an additional framework that can complement and be compared to previous studies. Specifically, we employ generalised additive models (GAMs; introduced by Hastie & Tibshirani, 1990) to expand theoretical chemical evolution models for the thin disc (in this case, those of Magrini et al. 2009) of the MW (which are characterised by widely spaced bins in age and radius) to investigate and differentiate the motion of thin disc stars (i.e. churn and blur/lack of interaction). Our GAM approach incorporates two independent variables consisting of chemical abundances ([Fe/H] or [Mg/H]; the analysis with [Mg/H] is in the Appendix) and $t$, the age of the Universe at the time the star was formed (roughly 13.8 Gyr [^1] We adopted the estimated age of the Universe as 13.8 Gyr (Planck Collaboration et al., 2020). minus the current estimated age of a star). The Galactocentric distance ($R$) serves as the response variable (i.e. the output of our regression model). Using GAM, we can effectively model the complex relationship between the chemical abundances and $t$ (independent variables) and $R$ without imposing rigid assumptions regarding the functional forms of these relationships.
@@
In contrast, the thick disc – despite its distinct [$\alpha$/Fe] enhancements, older stellar populations, and kinematically hot nature (see e.g. Soubiran et al., 2003; Bensby et al., 2011; Recio-Blanco et al., 2014; Bland-Hawthorn & Gerhard, 2016) – presents greater complexities that remain less well understood. Structures such as the halo and bulge pose even greater challenges for chemical evolution models, as their diverse formation mechanisms result in extreme variations in metallicity and dynamics. For instance, the halo reflects the earliest stages of Galactic chemical evolution and accretion events, being predominantly metal-poor (see e.g. Gonzalez-Jara2025). In contrast, the bulge, characterised by its old stellar populations and strikingly wide metallicity variations ($-1.5\lesssim\rm{[Fe/H]}\lesssim+0.5$; see Barbuy et al., 2018, for a comprehensive review), encapsulates the star formation history of the Galactic centre. Despite its importance, providing a self-consistent empirical chemo-dynamical model for the Galactic bulge remains a significant challenge due to its complexity (e.g. Grieco et al., 2012; Barbuy et al., 2018).
Insights from numerical simulations, particularly regarding metallicity gradients and dynamical evolution across these Galactic components, provide valuable context for understanding their formation and evolution (e.g. Grand et al., 2017; Tissera et al., 2022; Jara-Ferreira et al., 2024). However, the theoretical predictions must be compared with observations to ensure consistency with the MW’s history and evolution. Consequently, while the thin disc offers a tractable and data-rich framework for advancing our understanding of chemical evolution, expanding these models to incorporate other Galactic components – guided by both simulations and observational constraints – remains an essential and ongoing objective (see e.g. Minchev et al., 2013, for a study combining chemical enrichment models of the MW’s disc with numerical simulations of galactic discs).
This work forms the foundation of a broader investigation. Here, we focus on analysing results derived from the GAM and providing an in-depth exploration of our observed Gaia-ESO star sample. Our goal is to examine other aspects of these findings in future studies in order to better understand the causes and implications of the stellar motion trends revealed here.
This paper is organised as follows: Section 2 details the dataset and methodologies used in this study, thus providing the foundation for the subsequent analysis. The heart of the paper lies in Section 3, where we delve into the analysis and discuss the implications of our findings. Concluding the paper, Section 4 synthesises our insights and summarises the key takeaways and conclusions drawn from this phase of our investigation. Finally, we provide the details of our catalogue publicly available in the CDS in Section 5.
@@ 2.1 Data description
Figure 2: Scatter plot displaying the relationship between metallicity, [Fe/H], and stellar ages, $t_{\star}$, using unidam estimates for the entire sample. Data points are categorised by HC groups. On top, 1D-Gaussian kernel densities show age distribution; on the right, a similar view presents [Fe/H] distribution for each HC group. The window width for the kernel density estimation is computed using Scott’s method, with a default bandwidth adjustment of 1, which is the standard configuration of the seaborn package in python (Waskom, 2021); this configuration is consistently applied to all figures with Gaussian kernel densities throughout this paper. The data points and densities are colour-coded and stratified according to the different HC groups, similarly to Fig. 1.
@@
2.2 The Galactic chemical evolution models of Magrini et al.
@@
In Fig. 3 the radial profiles of the Magrini et al. (2009) models are shown. It is noticeable that the bottom panel, depicting [Fe/H], shows the smoothest curves (potentially the best estimated of the profiles), whereas all the other ones have constant or near-constant abundances after a certain $R$. This is an important feature to have in mind since this lack of smoothness can be detrimental to the implementation of the GAM. This is discussed further in Sect. 2.3. Moreover, there is a clear pattern of strong positive or negative correlations between the abundances in the models, as evidenced by the heatmap in Fig. 4.
@@
Figure 3: Chemical enrichment models of the MW as described by Magrini et al. (2009) illustrating the relative abundances of key elements in comparison to iron (Fe) across various Galactocentric radii ($R$). The elements, arranged sequentially by atomic number relative to Fe, include [O/Fe] at the top panel, followed by [Mg/Fe], [Si/Fe], [Ca/Fe], and culminating with [Fe/H] on the bottom panel. The colour gradation, ranging from a lively orange to black, represents the ages of the Universe at the time of star formation within the MW, starting at $t$=2.1 Gyr and extending to $t$=13.7 Gyr. This visual spectrum effectively portrays the chronological progression of the Galaxy’s chemical composition. The legend, for clearer visualisation, is incorporated within the [Ca/Fe] subplot.
Figure 4: Heatmap visualisation of the correlation matrix showing the pairwise correlation coefficients between elemental abundances relative to iron ([O/Fe], [Mg/Fe], [Si/Fe], [Ca/Fe]) and the iron abundance relative to hydrogen ([Fe/H]). Orange indicates a strong positive correlation, while brown signifies a strong negative correlation, demonstrating the tight coupling between these elemental abundances in the context of Magrini et al.’s model. The correlation matrix was estimated via Spearman’s rank correlation coefficient (Spearman, 1904).
@@ 2.3 Generalised additive models
2.3.1 Why we adopt generalised additive models
@@
Generalised additive models (introduced by Hastie & Tibshirani, 1990), an extension of GLMs, offer even greater flexibility by allowing for non-linear relationships between predictors and the response variable, using smooth functions. Despite their potential, to our knowledge, GAMs have been previously used fewer times in astrophysics, such as in photometric redshift estimation (Beck et al., 2017), in the analysis of ionising photon escape fractions from dark haloes (Hattab et al., 2019), and more recently in the examination of the relationship between galaxy structure and star formation rate (Stephenson et al., 2024).
In recent years, astrophysics seems to have skipped over these ‘intermediate’ techniques, with many researchers jumping directly from classical methods to highly complex approaches such as Gaussian processes and deep learning. While these advanced techniques are undoubtedly valuable, methods such as GLMs and GAMs offer elegant and interpretable solutions that are more straightforward to implement and understand. This accessibility makes them highly valuable in a range of disciplines, shown by its use in ecology (e.g. Simpson, 2018; Pedersen et al., 2019; Kosicki, 2020), linguistics (e.g. Wieling et al., 2016; Tomaschek et al., 2018), medicine (e.g. de Souza & Berger, 2021), to mention a few.
2.3.2 The general form of a generalised additive model